### v1

@@ 1 Introduction
One of the key challenges in healthcare engineering is the automatic analysis of ECG signals for early detection of cardiac abnormalities and timely clinical intervention. Despite digital recording, interpretation of 12‐lead ECGs still depends on expert visual assessment of waveforms and remains challenging due to inter‐individual variability in gender, electrode placement, and electrophysiological properties [30]. Conversely, recent innovations in healthcare have made the ECG increasingly a commodity, while training in ECG interpretation remains limited, placing a greater workload on experts and potentially hampering the early detection of cardiac patients [18].
@@
On the other hand, the availability of large‐scale, clinically accurate 12‐lead ECG datasets (e.g., PTB‐XL [33, 32] and PTB‐XL+ [26]) has enabled the development of high‐performance deep learning classifiers capable of detecting multiple cardiac conditions within a single recording. However, the inherent complexity of these models renders their decision‐making processes opaque to end users [12], underscoring the need for tailored explainable AI techniques in ECG interpretation.
@@
Despite progress in SHAP-based local explanations such as WindowSHAP [16], current methods might be too slow for real-time deployment [31]. Counterfactual generation techniques frequently violate ECG morphology and temporal coherence, yielding physiologically implausible waveforms.  [3] describe a lightweight ECG analysis system tailored for wearable and embedded platforms. They highlight the demand for efficient models capable of real-time analysis in resource-limited settings, yet many XAI techniques still introduce substantial computational overhead. This persistent shortfall in explainability performance motivates our work: our framework delivers potentially low-latency inference, actionable counterfactual explanations, and the potential for point-of-care deployment on standard clinical workstations.
@@
In recent works, explainable AI methods have been applied to ECG interpretation with promising results. Image-based techniques, such as those of [1], overlay heatmaps on 12-lead ECG images to highlight diagnostically relevant waveform features and improve model interpretability. They further advocate for advanced XAI strategies: counterfactual explanations and rule-based methods, to boost both transparency and robustness. Similarly, [28] employed DiCE-based counterfactuals [14] on the PTB-XL dataset, validating generated examples with two cardiologists and demonstrating their clinical relevance. Building on these complementary approaches, our framework combines rule-based explanations, heatmaps and efficient counterfactual generation to deliver real-time, actionable insights.
@@
In the context of ECG counterfactual explanations, sparsity refers to the proportion of the signal that requires modification to achieve a class transition. A sparse counterfactual modifies only a minimal subset of time points across the 12 leads, rather than perturbing the entire recording. This sparsity is crucial for clinical interpretability since sparse explanations highlight specific waveform segments that drive diagnostic changes instead of presenting wholesale signal replacements. By focusing clinical attention on these critical regions, which states a challenge in the literature, sparse counterfactuals align with how cardiologists naturally interpret ECGs—examining specific morphological features rather than global waveform differences.
@@
Our framework addresses these challenges by combining SHAP-derived rule extraction and prototype-based sparse counterfactual generation. First, we compute SHAP values and select features above a global threshold, forming concise interval rules. Next, we measure distances between rule-defined series using Dynamic Time Warping (DTW) [22] and employ clustering followed by medoid selection to extract representative prototypes for each class. Finally, we apply our optimisation step, which consists of aligning and scaling these prototypes to the query ECG by matching characteristic R-peaks. To enrich comprehensibility, we interviewed clinicians, including cardiology and electrophysiology experts, using semi-structured, one-on-one sessions. Participants reviewed lead-wise importance charts, signal overlays, and both single- and multi-overlay counterfactual examples annotated with interval rules. They rated readability, design clarity, and information density, and offered improvement suggestions.
@@
In summary, our method delivers fast, physiologically valid counterfactual ECG explanations. It combines SHAP-derived interval rules with precomputed prototypes to preserve both waveform morphology and global rhythm. By precomputing prototypes, counterfactuals are generated in near real time with the cost of querying and optimization. Clinician interviews confirmed the readability, clarity, and appropriate information density of our visualizations. This approach strengthens transparency and trust in AI-driven ECG analysis.
@@
The remainder of this article is structured as follows. Section 2 reviews the background and related work. Section 3 describes our proposed method for generating counterfactual explanations of ECG signals, including adaptation of the Post-hoc Attribution Rule extraction (PHAR) framework (Section 3.1), prototype selection (Section 3.2), and the explanation generation process (Section 3.3). Section 4 presents the experimental evaluation, including quantitative results and qualitative interviews with ECG analysts. Section 5 discusses the advantages and limitations of our approach and outlines directions for future work. Finally, Section 6 concludes the article by summarizing the main contributions. Additional details are provided in the appendices: Appendix A describes the details of ECG model training and calibration, Appendix B reports the metrics for ECG counterfactual evaluation, and Appendix C presents example ECG counterfactual visualizations from the proposed method.
@@ 2 Background
2.1 XAI in Healthcare
Recent work by [20] emphasizes the growing adoption of AI and XAI systems in healthcare. In the study’s semi‑structured interviews, clinicians identified feature‑importance scores as among the most immediately interpretable explanation methods. These visualizations, often simple bar charts, enable practitioners to quickly recognize which patient characteristics most influence the model’s prediction without requiring understanding of complex algorithmic techniques such as SHAP or LIME. However, clinicians also cautioned that this simplicity can be misleading. Some remarked that interpreting these scores may feel deceptively straightforward, raising concerns about oversimplification. To establish trust, they emphasized the need for additional contextual information, such as uncertainty estimates or reference distributions, to support these explanations [21].
While counterfactual and rule‐based explanations offer deeper contextual understanding, they demand additional time and cognitive effort to assess. Visual representations, such as bar charts and signal overlays, are essential for translating model outputs into comprehensible insights at the point of care. Successful deployment depends on solutions that integrate seamlessly into existing workflows and on multidisciplinary collaboration between developers, researchers, and healthcare professionals.
@@ 2.1 XAI in Healthcare
Holzinger et al. [11] argue that effective XAI in medicine requires intuitive explanation interfaces and tight human‐in‐the‐loop integration to ensure clinical usability. They emphasize the trade‐off between ante‐hoc models, which offer built‐in transparency but may sacrifice accuracy, and post‐hoc methods that provide flexible explanations for black‐box predictors. Moreover, they advocate for causal frameworks capable of generating counterfactual insights, thereby aligning technical explainability with clinicians’ causal reasoning.
@@
[23] review 213 studies on XAI in cardiology and report that 43% of works do not validate explanations, 37% rely solely on literature comparison, 11% involve clinical experts, and 11% apply statistical or proxy-grounded evaluation. Commonly used techniques for ECG interpretation include SHAP (global and local importance scores) and Grad-CAM (activation maps), sometimes augmented with LIME. In the minority of studies engaging cardiologists, evaluation focuses on the alignment between model-highlighted critical points and expert interpretation. The authors emphasize the necessity of application-grounded validation with cardiologists, the adoption of proxy-grounded methods (which assess explanation fidelity by removing attributed features and retraining the model), permutation tests (which quantify the effect of feature shuffling on predictions) and a hybrid evaluation strategy combining human-, proxy-, and literature-grounded approaches to ensure clinical relevance and trustworthiness. A similar comprehensive evaluation approach is adopted by [7] and [6], who employ addition and deletion metrics that systematically assess explanation fidelity by removing or adding attributed features and measuring the resulting impact on model predictions. These proxy-grounded methods provide quantitative measures of explanation quality by evaluating how feature modifications affect classification performance, complementing human-centered evaluation approaches. Such multi-faceted evaluation frameworks are essential for establishing the reliability and clinical utility of XAI systems in healthcare applications.
@@ 2.2 Interactive Visualization for Expert Feedback
Interactive machine learning (iML) highlights the critical role of end‐user engagement through iterative model refinement. Wondimu et al. [36] demonstrate that interactive visualizations, such as multi‐view decision path diagrams and activation heatmaps, empower users to inspect model behavior, adjust parameters in real time, and directly observe the impact of their inputs. The authors emphasize that transparency, user control, and ongoing human–machine collaboration are essential foundations for trustworthy AI in high‐stakes domains.
Theissler et al. [29] propose a four‐fold taxonomy of XAI for time‐series: time‐point attribution, subsequence methods, instance‐based strategies (including counterfactuals), and rule/logic frameworks. They observe that ECG applications predominantly rely on attribution heatmaps and subsequent highlighting, yet often lack interactive evaluation and standardized benchmarks. The authors emphasize visualization (saliency maps, activation heatmaps, annotated overlays) as essential for translating model decisions into experts’ insights. Key research gaps include the development of higher‐order and truly model‐agnostic explanations, domain‐specific interpretability techniques, user‐friendly interfaces, and advanced evaluation protocols.
@@ 2.3 Deep Learning for ECG Classification
Deep neural networks achieve high accuracy on multivariate time series tasks [34]. Classification models have shown state-of-the-art performance. An exemplary deep learning model for ECG analysis is described by [10], which achieves cardiologist‐level arrhythmia detection but does not incorporate any explainability (XAI) techniques.  [3] develop lightweight 1D CNN models for on‐device arrhythmia classification and integrate popular XAI methods: Grad‐CAM to produce local activation maps that highlight key signal components (e.g., QRS complexes, T waves) and SHAP to assign per‐timepoint attribution scores across different diagnoses. While these well‐established techniques yield valuable model insights, their raw visualizations can be cluttered and insufficiently streamlined for fast clinical interpretation.
@@ 2.4 Counterfactual Explanations
[8] introduces Native Guide, an instance‐based method for obtaining counterfactual explanations in time‐series classification. The approach retrieves the nearest opposite‐class exemplar from a reference database and adapts its key subsequences using classifier‐derived weight vectors, yielding counterfactuals that align closely with the original data distribution. This exemplar search can be computationally demanding. The authors validate their method on multiple UCR benchmarks, including ECG200111https://www.timeseriesclassification.com/description.php?Dataset=ECG200 [17], which contains pre‐segmented single‐beat ECG signals. These simplified datasets are less complex than real‐world clinical recordings. They report superior performance over gradient‐based counterfactual techniques in terms of proximity, sparsity, plausibility, and diversity.

### v2

@@ 1 Introduction
One of the key challenges in healthcare engineering is the automatic analysis of ECG signals for early detection of cardiac abnormalities and timely clinical intervention. Despite digital recording, interpretation of 12‐lead ECGs still depends on expert visual assessment of waveforms and remains challenging due to inter‐individual variability in gender, electrode placement, and electrophysiological properties [VANDAM202167]. Conversely, recent innovations in healthcare have made the ECG increasingly a commodity, while training in ECG interpretation remains limited, placing a greater workload on experts and potentially hampering the early detection of cardiac patients [PRONIEWSKA202345].
@@
On the other hand, the availability of large‐scale, clinically accurate 12‐lead ECG datasets (e.g., PTB‐XL [wagner_2020_ptbxl, wagner_2022_ptbxl] and PTB‐XL+ [Strodthoff2023] has enabled the development of high‐performance deep learning classifiers capable of detecting multiple cardiac conditions within a single recording. However, the inherent complexity of these models renders their decision‐making processes opaque to end users [Holzinger2019], underscoring the need for tailored explainable AI techniques in ECG interpretation. Crucially, the challenge extends beyond the algorithm itself. As [SCHMITT2023100520] argues about safety-critical digital ecosystems, effective deployment requires seamless integration, real-time responsiveness, and strict governance. Our framework aligns with these principles by providing efficient and verifiable explanations. Those requirements are equally vital for clinical decision-support systems. Consequently, there is a pressing need for tailored explainable AI techniques that not only clarify model outputs but also fit within rigorous healthcare workflows.
@@
Despite progress in SHAP-based local explanations [Lundberg2017] such as WindowSHAP [WindowSHAP2023], current methods might be too slow for real-time deployment [vanlooveren2020interpretablecounterfactualexplanationsguided]. Counterfactual generation techniques frequently violate ECG morphology and temporal coherence, yielding physiologically implausible waveforms.  [baig2025arrhythmiavisionresourceconsciousdeeplearning] describe a lightweight ECG analysis system tailored for wearable and embedded platforms. They highlight the demand for efficient models capable of real-time analysis in resource-limited settings, yet many XAI techniques still introduce substantial computational overhead. This persistent shortfall in explainability performance motivates our work: our framework delivers potentially low-latency inference, actionable counterfactual explanations, which are physiologically faithful as they are grounded in real patient prototypes, and the potential for point-of-care deployment on standard clinical workstations.
@@
In recent works, explainable AI methods have been applied to ECG interpretation with promising results. Image-based techniques, such as those of [image_based_ecg_2023], overlay heatmaps on 12-lead ECG images to highlight diagnostically relevant waveform features and improve model interpretability. They further advocate for advanced XAI strategies: counterfactual explanations and rule-based methods, to boost both transparency and robustness. Similarly, [TANYEL2025107227] employed DiCE-based counterfactuals [mothilal2020dice] on the PTB-XL dataset, validating generated examples with two cardiologists and demonstrating their clinical relevance. Building on these complementary approaches, our framework combines rule-based explanations, heatmaps and efficient counterfactual generation to deliver real-time, actionable insights.
@@
In the context of ECG counterfactual explanations, sparsity refers to the proportion of the signal that requires modification to achieve a class transition. A sparse counterfactual modifies only a minimal subset of time points across the 12 leads, rather than perturbing the entire recording. This sparsity is crucial for clinical interpretability since sparse explanations highlight specific waveform segments that drive diagnostic changes instead of presenting wholesale signal replacements. By focusing clinical attention on these critical regions, which states a challenge in the literature, sparse counterfactuals align with how cardiologists naturally interpret ECGs, examining specific morphological features rather than global waveform differences.
Figure 1: Overview of the proposed framework. ECG data is classified by an AI model, and SHAP-based rule extraction identifies important features. Rules are clustered per class to select prototype ECGs. For a query, the system retrieves prototypes, applies R-peak alignment and sparsity optimization, and generates counterfactual explanations.
@@
Our framework addresses these challenges by combining SHAP-derived rule extraction and prototype-based sparse counterfactual generation (see Figure 1 for an overview). First, employing the Post-hoc Attribution Rule extraction (PHAR) framework [mozolewski2025phar], we compute SHAP values and select features above a global threshold, forming concise interval rules. Next, we measure distances between rule-defined series using Dynamic Time Warping (DTW) [DWT1163055] and employ clustering followed by medoid selection to extract representative prototypes for each class. Finally, we apply our optimisation step, which consists of aligning and scaling these prototypes to the query ECG by matching characteristic R-peaks. To enrich comprehensibility, we interviewed clinicians, including cardiology and electrophysiology experts, using semi-structured, one-on-one sessions. Participants reviewed lead-wise importance charts, signal overlays, and both single- and multi-overlay counterfactual examples annotated with interval rules. They rated readability, design clarity, and information density, and offered improvement suggestions.
@@
In summary, our method delivers fast, physiologically valid counterfactual ECG explanations. It combines SHAP-derived interval rules with precomputed prototypes to preserve both waveform morphology and global rhythm. By precomputing prototypes, counterfactuals are generated in near real time with the cost of querying and optimization. Preliminary clinician interviews provided initial positive feedback on the readability, clarity, and information density of our visualizations. This approach strengthens transparency and trust in AI-driven ECG analysis.
@@
The remainder of this article is structured as follows. Section 2 reviews the background and related work. Section 3 describes our proposed method for generating counterfactual explanations of ECG signals, including adaptation of the Post-hoc Attribution Rule extraction (PHAR) framework (Section 3.1), prototype selection (Section 3.2), and the explanation generation process (Section 3.3). Section 4 presents the experimental evaluation, including quantitative results and qualitative interviews with ECG analysts. Section 5 discusses the advantages and limitations of our approach and outlines directions for future work. Finally, Section 6 concludes the article by summarizing the main contributions. Additional details are provided in the appendices: Appendix A describes the details of ECG model training and calibration, Appendix B reports the metrics for ECG counterfactual evaluation, Appendix C presents example ECG counterfactual visualizations from the proposed method, and Appendix D outlines the expert evaluation protocol.
@@ 2 Background
2.1 Explainable AI in Healthcare
Recent work by [cliniciansvoice_2025] emphasizes the growing adoption of AI and XAI systems in healthcare. In the study’s semi-structured interviews, clinicians prioritized feature-importance scores due to their immediate visual interpretability, which allows practitioners to identify key predictive factors without requiring deep knowledge of underlying algorithms like SHAP [Lundberg2017] or LIME [LIME]. However, respondents cautioned that this abstract simplicity can be deceptive, provoking skepticism regarding the validity of reducing complex clinical scenarios to basic additive factors.
In the specialized domain of clinical decision support, SHAP and LIME have emerged as the most prominent paradigms for interpreting complex medical predictions [Sannidhi2022AutonomousDiseasePredictions, Ahmed2025DiabetesXAI]. A comprehensive literature review conducted by [Manimaran2025ECGLiteratureReview] reveals that SHAP is the most frequent choice for electrocardiogram (ECG) analytics, appearing in 21.9% of studies while LIME follows at 9.4% [Manimaran2025ECGLiteratureReview]. While both frameworks facilitate transparency, they utilize fundamentally different mechanisms since SHAP is rooted in additive feature attribution from game theory and LIME focuses on local feature perturbation [Salih2025ShapLime]. However, as highlighted by Salih et al., post-hoc attribution methods like SHAP and LIME often struggle with highly correlated biomedical signals, which may obscure true physiological relationships. This inherent instability, as noted by [Singh2025Beyond], poses a significant risk for misleading clinical interpretations if raw importance scores are used in isolation.
As demonstrated in the clinical interviews conducted by [cliniciansvoice_2025], alternative paradigms such as counterfactual explanations and rule-based models were acknowledged for offering deeper insights or mirroring human reasoning, but faced practical criticism. Specifically, counterfactuals were viewed as potentially causally inconsistent or irrelevant due to their focus on hypothetical patients, whereas rule-based systems were deemed too cumbersome for rapid decision-making.
@@ 2.1 Explainable AI in Healthcare
Holzinger et al. [Holzinger2019Causability] argue that effective XAI in medicine requires intuitive explanation interfaces and tight human‐in‐the‐loop integration to ensure clinical usability. They emphasize the trade‐off between ante‐hoc models, which offer built‐in transparency but may sacrifice accuracy, and post‐hoc methods that provide flexible explanations for black‐box predictors. Moreover, they advocate for causal frameworks capable of generating counterfactual insights, thereby aligning technical explainability with clinicians’ reasoning.
The relevance of XAI in clinical decision-making is further underscored by a systematic review by [Antoniadi2021Systematic], which identifies that the core value of interpretability in Clinical Decision Support Systems (CDSS) is to provide local explanations that support the decision-making process of medical practitioners’ in specific cases. They observe that XAI research is mainly focused on tabular datasets, while other modalities are less frequent. Crucially, the authors highlight that ensuring adoption, as opposed to simply technical acceptance, requires bridging the lack of user-centric validation through expert-engaged studies.
@@
This ongoing challenge of validation is echoed by [Salih2024EvalXAI], who review 213 studies on XAI in cardiology and report that 43% of works do not validate explanations, 37% rely solely on literature comparison, 11% involve clinical experts, and 11% apply statistical or proxy-grounded evaluation. Commonly used techniques for ECG interpretation include SHAP (global and local importance scores) and Grad-CAM (activation maps), sometimes augmented with LIME. In the minority of studies engaging cardiologists, evaluation focuses on the alignment between model-highlighted critical points and expert interpretation. The authors emphasize the necessity of application-grounded validation with cardiologists, the adoption of proxy-grounded methods (which assess explanation fidelity by removing attributed features and retraining the model), permutation tests (which quantify the effect of feature shuffling on predictions) and a hybrid evaluation strategy combining human-, proxy-, and literature-grounded approaches to ensure clinical relevance and technical reliability.
A similar comprehensive evaluation approach is adopted by [darias2024empirical] and [darias2025evaluating], who employ addition and deletion metrics that systematically assess explanation fidelity by removing or adding attributed features and measuring the resulting impact on model predictions. These proxy-grounded methods provide quantitative measures of explanation quality by evaluating how feature modifications affect classification performance, complementing human-centered evaluation approaches. Such multi-faceted evaluation frameworks are essential for establishing the reliability and clinical utility of XAI systems in healthcare applications.
The necessity of such multi-faceted frameworks is further underscored by [Amann2022ToExplain], who argue that the value of explainability in CDSS is inherently context-dependent. They emphasize that XAI should not merely provide data, but facilitate “meaning-making” for the clinician by balancing technical feasibility with the user’s cognitive constraints. This perspective provides a principled foundation for our methodology. Specifically, our sparsity optimization directly addresses the requirement to minimize information overload, ensuring that only the most critical diagnostic changes are presented. Furthermore, the use of R-peak alignment and regional highlighting ensures that the explanations align with clinical intuition and facilitate rapid interpretation. By adhering to these design principles, our system moves beyond raw feature attribution toward a task-oriented CDSS that supports actionable decision-making.
@@ 2.2 Interactive Visualization for Expert Feedback
Interactive machine learning (iML) highlights the critical role of end‐user engagement through iterative model refinement. Wondimu et al. [Wondimu2022InteractiveML] demonstrate that interactive visualizations, such as multi‐view decision path diagrams and activation heatmaps, empower users to inspect model behavior, adjust parameters in real time, and directly observe the impact of their inputs. The authors emphasize that transparency, user control, and ongoing human–machine collaboration are essential foundations for trustworthy AI in high‐stakes domains.
Theissler et al. [xai_ts_class_2022] propose a four‐fold taxonomy of XAI for time‐series: time‐point attribution, subsequence methods, instance‐based strategies (including counterfactuals), and rule/logic frameworks. They observe that ECG applications predominantly rely on attribution heatmaps and subsequent highlighting, yet often lack interactive evaluation and standardized benchmarks. The authors emphasize visualization (saliency maps, activation heatmaps, annotated overlays) as essential for translating model decisions into experts’ insights. Key research gaps include the development of higher‐order and truly model‐agnostic explanations, domain‐specific interpretability techniques, user‐friendly interfaces, and advanced evaluation protocols.
@@ 2.3 Deep Learning for ECG Classification
Deep neural networks achieve high accuracy on multivariate time series tasks [Wang2024]. Classification models have shown state-of-the-art performance. An exemplary deep learning model for ECG analysis is described by [Hannun2019], which achieves cardiologist‐level arrhythmia detection but does not incorporate any explainability (XAI) techniques.  [baig2025arrhythmiavisionresourceconsciousdeeplearning] develop lightweight 1D CNN models for on‐device arrhythmia classification and integrate popular XAI methods: Grad‐CAM to produce local activation maps that highlight key signal components (e.g., QRS complexes, T waves) and SHAP to assign per‐timepoint attribution scores across different diagnoses.
Furthermore, beyond standard CNN architectures, deep generative models have become integral to physiological signal analysis. A recent systematic review by [NEIFAR2025103127] confirms their extensive application in ECG processing, identifying Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) as the predominant architectures, primarily used for data augmentation and signal denoising. However, in the context of resource-constrained environments, flow-based approaches offer distinct advantages. [Ibrahim2025EnhancingEC] conducted a comprehensive comparison of unsupervised anomaly detection filters, including Deep SVDD, Autoencoders, VAEs, and diffusion models. Their study demonstrated that Normalizing Flows (NF) achieve a superior trade-off between computational efficiency and detection performance, often outperforming reconstruction-based approaches. This capacity to accurately model complex probability densities makes NF a compelling choice for robust feature extraction in lightweight clinical systems.
@@ 2.4 Counterfactual Explanations
While standard CNN classifiers and attribution-based XAI techniques yield valuable insights, their raw visualizations can be cluttered and insufficiently streamlined for fast clinical interpretation.
@@
To overcome the visual complexity of raw attribution maps, instance-based explanations provide intuitive alternatives by showing how an input would need to change to achieve a different outcome.  [Delaney2021InstanceBasedCounterfactual] introduces Native Guide, an instance‐based method for obtaining counterfactual explanations in time‐series classification. The approach retrieves the nearest opposite‐class exemplar from a reference database and adapts its key subsequences using classifier‐derived weight vectors, yielding counterfactuals that align closely with the original data distribution. The authors validate their method on multiple UCR benchmarks, including ECG200111https://www.timeseriesclassification.com/description.php?Dataset=ECG200 [Olszewski2001], which contains pre‐segmented single‐beat electrocardiogram (ECG) signals. These simplified datasets are less complex than real‐world clinical recordings. They report superior performance over gradient‐based counterfactual techniques in terms of proximity, sparsity, plausibility, and diversity.
Work by [Karlsson2020Tweaking] formalized the problem of “time series tweaking”, distinguishing between local strategies (modifying specific shapelets) and global strategies (using nearest neighbors). In the local setting, they leverage Random Shapelet Forests (RSF) to detect discriminative subsequences, subsequently modifying the query signal to match the shapelet values required for a class change. Conversely, their global strategy utilizes a $k$-Nearest Neighbor ($k$-NN) approach to retrieve an existing instance from the target class, positing that historical examples offer inherently plausible explanations. While their global $k$-NN approach effectively identifies plausible counterfactuals in general time-series domains, its direct application to ECG analysis is limited by the lack of temporal alignment, as phase shifts (e.g., R-peak offsets) can distort the morphological comparison essential for cardiac diagnosis.
Positioning their work against these instance-based baselines, specifically comparing performance with [Delaney2021InstanceBasedCounterfactual] and the shapelet approaches of [Karlsson2020Tweaking][Wang2024Glacier] proposed Glacier, a framework generating counterfactuals via gradient search within either the original feature space or an auto-encoder’s latent representation. They report superior proximity and compactness compared to the aforementioned retrieval methods. A central contribution of their work is the integration of “local constraints” to guide the gradient descent, aiming to restrict modifications to specific time segments and suppress high-frequency noise. However, despite these regularizations, the method remains fundamentally reliant on numerical perturbation. Such mathematical deformation, while optimizing for decision boundaries, lacks explicit biological constraints, risking the introduction of signal artifacts that may not align with cardiac electrophysiology.
The potential of counterfactuals for clinical decision support has been demonstrated in the context of ECG analysis for datasets more closely related to clinical practice.   [TANYEL2025107227] introduced the VCCE method for the detection of myocardial infarction (MI) using the PTB-XL dataset. By utilizing the DiCE framework, they generate counterfactual clues through feature perturbations, providing clinicians with visual evidence of diagnostic transitions. While their work highlights the high clinical alignment of optimization-based counterfactuals in binary classification, such approaches can be computationally intensive and may lack explicit rhythm consistency. In contrast, our framework extends beyond binary MI detection to a multi-class setting and addresses the need for physiologically-consistent rhythm alignment and real-time inference.
While counterfactual and rule-based explanations offer deeper contextual understanding, they demand additional time and cognitive effort to assess. Successful deployment depends on so-called unremarkable AI solutions that integrate seamlessly into existing workflows [Yang2019Unremarkable] and on multidisciplinary collaboration between developers, researchers, and healthcare professionals.
2.5 Prototype-based Explanations